# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ 4.2 Terminating Algorithm
The subroutine ApproxCount. We first introduce the subroutine ApproxCount, whose formal description and proof of correctness are found in Appendix A. The purpose of ApproxCount is to compute an approximation $n^{\prime}$ of the total number of processes $n$ (or report various types of failure). It takes as input a view $\mathcal{V}$ of a process, the number of leaders $\ell$, and two integer parameters $s$ and $x$, representing the index of a level of $\mathcal{V}$ and the anonymity of a leader node in $L_{s}$, respectively.
@@
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\tau$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\tau$ does not exist, the procedure immediately returns the error code $n^{\prime}=-1$). We define the discrepancy $\delta$ as the ratio $x/a(\tau)$. Clearly, $\delta\leq\ell$. Note that, since $a(\tau)$ is not a-priori known by the process executing ApproxCount, then neither is $\delta$.
@@
Conditional anonymities. ApproxCount starts by assuming that the anonymity of $\tau$ is $x$, and makes deductions on other anonymities based on this assumption. Thus, we will distinguish between the actual anonymity of a node $a(v)$ and the conditional anonymity $a^{\prime}(v)=\delta a(v)$ that ApproxCount may compute under the initial assumption that $a^{\prime}(\tau)=x=\delta a(\tau)$.
@@
Overview of ApproxCount. The procedure ApproxCount scans the levels of $\mathcal{V}$ starting from $L_{s}$, making “guesses” on the conditional anonymities of nodes based on already known conditional anonymities. Generalizing some lemmas from [24], we develop a criterion to determine when a guess is correct. This yields more nodes with known conditional anonymities, and therefore more guesses (the details are in Appendix A). As soon as it has obtained enough information, the procedure stops and returns $(n^{\prime},t)$, where $L_{t}$ is the level scanned thus far. If the information gathered satisfies certain criteria, then $n^{\prime}$ is an approximation of $n$. Otherwise, $n^{\prime}$ is an error code, as detailed below.
@@
Error codes. If $L_{s}$ contains no leader nodes, the procedure returns the error code $n^{\prime}=-1$. If, before gathering enough information on $n$, the procedure encounters a descendant of $\tau$ with more than one child in $\mathcal{V}$, it returns the error code $n^{\prime}=-2$. If it determines that the conditional anonymity of a node is not an integer, it returns the error code $n^{\prime}=-3$. Finally, if it determines that the sum $\ell^{\prime}$ of the conditional anonymities of the leader nodes is not $\ell$, it returns $n^{\prime}=-1$ if $\ell^{\prime}<\ell$ and $n^{\prime}=-3$ if $\ell^{\prime}>\ell$.
@@
Correctness of ApproxCount. The following lemma gives some conditions that guarantee that ApproxCount has the expected behavior; it also gives some bounds on the number of rounds it takes for ApproxCount to produce an approximation $n^{\prime}$ of $n$, as well as a criterion to determine if $n^{\prime}=n$. The lemma’s proof is rather lengthy and technical, and is deferred to Appendix A.
@@
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\tau$ exists and $x\geq a(\tau)$. Let $\tau^{\prime}$ be the (unique) descendant of $\tau$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@
    If $x=a(\tau)=a(\tau^{\prime})$, then $n^{\prime}\neq-3$.
@@
    If $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$ and $a(\tau)=a(\tau^{\prime})$, then $n^{\prime}=n$.
@@
    If $t^{\prime}\geq s+(\ell+2)n-1$, then $s\leq t\leq s+(\ell+1)n-1$ and $n^{\prime}\neq-1$. Moreover, if $n^{\prime}=-2$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.∎
Our terminating algorithm assumes that all processes know the number of leaders $\ell\geq 1$ and the dynamic disconnectivity $T$. Again, this is justified by Proposition 5.5 and Proposition 2.2.
@@
There is an algorithm that computes $F_{GC}$ in all $T$-union-connected anonymous networks with $\ell\geq 1$ leaders and terminates in at most $(\ell^{2}+\ell+1)Tn$ rounds, assuming that $\ell$ and $T$ are known to all processes, but assuming no knowledge of $n$.
@@
Due to Proposition 2.3, since $T$ is known and appears as a factor in the claimed running time, we can assume that $T=1$ without loss of generality. Also, note that determining $n$ is enough to compute $F_{GC}$. Indeed, if a process determines $n$ at round $t^{\prime}$, it can wait until round $\max\{t^{\prime},2Tn\}$ and run the algorithm in Theorem 4.1, which is guaranteed to give the correct output by that time.
In order to determine $n$ assuming that $T=1$, we let each process run the algorithm in Listing 2 with input $(\mathcal{V},\ell)$, where $\mathcal{V}$ is the view of the process at the current round $t^{\prime}$. We will prove that this algorithm returns a positive integer (as opposed to “Unknown”) within $(\ell^{2}+\ell+1)n$ rounds, and the returned number is indeed the correct size of the system $n$.
@@ 8   For $x:=\ell$ downto $1$
9      Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$      # see Listing 3 in Appendix A
@@ 17   Assign $s:=t^{\ast}+1$
In order to compute an approximation of $n$, say in an interval of levels $\mathcal{L}_{i}$ starting at $L_{s}$, the algorithm goes through at most $\ell$ phases (Lines 8–16). The first phase begins by calling ApproxCount with starting level $L_{s}$ and $x=\ell$, i.e., the maximum possible value for the anonymity of a leader node (Line 9). Specifically, ApproxCount chooses a leader node in $\tau\in L_{s}$ and tries to estimate $n$ using as few levels as possible.
Let $(n^{\prime},t)$ be the pair of values returned by ApproxCount. If $n^{\prime}=-1$, this is evidence that $\mathcal{V}$ is still missing some relevant nodes, and therefore “Unknown” is immediately returned (Line 11). If $n^{\prime}=-2$, then a descendant of $\tau$ with multiple children in $\mathcal{V}$ was found, say at level $L_{t}$, before an approximation of $n$ could be determined. As this is an undesirable event, the algorithm moves $\mathcal{L}_{i}$ after $L_{t}$ and tries again to estimate $n$ (Line 12). If $n^{\prime}=-3$, then $x$ may not be the correct anonymity of the leader node $\tau$ (see the description of ApproxCount), and therefore the algorithm calls ApproxCount again, with the same starting level $L_{s}$, but now with $x=\ell-1$. If $n^{\prime}=-3$ is returned again, then $x=\ell-2$ is tried, and so on. After all possible assignments down to $x=1$ have failed, the algorithm just moves $\mathcal{L}_{i}$ forward and tries again from $x=\ell$.
@@
Finally, when $\ell-b+1$ approximations of $n$ (all equal to $n^{\ast}$) have been found, a correctness check is performed: the algorithm takes the last level $L_{t^{\ast}}$ visited thus far; if the current round $t^{\prime}$ satisfies $t^{\prime}\geq t^{\ast}+n^{\ast}$, then $n^{\ast}$ is accepted as correct; otherwise “Unknown” is returned (Lines 18–20).
@@
Correctness and running time. We will prove that, if the output of Listing 2 is not “Unknown”, then it is indeed the number of processes, i.e., $n^{\ast}=n$. Since the $\ell-b+1$ approximations of $n$ have been computed on disjoint intervals of levels, there is at least one such interval, say $\mathcal{L}_{j}$, where no leader node in the history tree has more than one child (because there can be at most $\ell$ leader branches). With the notation of Lemma 4.2, this implies that $a(\tau)=a(\tau^{\prime})$ whenever ApproxCount is called in $\mathcal{L}_{j}$. Also, since the option $x=\ell$ is tried first, the assumption $x\geq a(\tau)$ of Lemma 4.2 is initially satisfied. Note that ApproxCount cannot return $n^{\prime}=-1$ or $n^{\prime}=-2$, or else $\mathcal{L}_{j}$ would not yield any approximation of $n$. Moreover, by statement (ii) and by the terminating condition (Line 19), if $n^{\prime}>0$ while $x\geq a(\tau)$, then $n^{\ast}=n^{\prime}=n$. On the other hand, by statement (i), we necessarily have $n^{\prime}>0$ by the time $x=a(\tau)$.
@@
It remains to prove that Listing 2 actually gives an output other than “Unknown” within the claimed number of rounds; it suffices to show that it does so if it is executed at round $t^{\prime}=(\ell^{2}+\ell+1)n$. It is known that all nodes in the first $t^{\prime}-n=\ell(\ell+1)n$ levels of the history tree are contained in the view $\mathcal{V}$ at round $t^{\prime}$ (cf. [24, Corollary 4.3]). Also, it is straightforward to prove by induction that the assumption of statement (iii) of Lemma 4.2 holds every time ApproxCount is invoked. Indeed, in any interval of $(\ell+1)n$ levels, either a branching leader node is found or a new approximation of $n$ is computed. Since there can be at most $\ell$ leader branches, at least one approximation of $n$ is computed within $\ell(\ell+1)n$ levels. Because all nodes in these levels must appear in $\mathcal{V}$, the condition $a(\tau)=a(\tau^{\prime})$ of Lemma 4.2 is satisfied in all intervals $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, …, $\mathcal{L}_{\ell-b+1}$. Reasoning as in the previous paragraph, we conclude that all such intervals must yield the correct approximation of $n$. So, every time Line 15 is executed, we have $n^{\ast}=n^{\prime}$, and the algorithm cannot return “Unknown”. ∎
5 Negative Results
In this section we provide simple proofs of several negative results and counterexamples, some of which are well known (in particular, Proposition 5.1 is implied by [28, Theorem III.1]). The purpose is to justify all of the assumptions made in Sections 3 and 4.
5.1 Leaderless Networks
Proposition 5.1.
No function other than the frequency-based multi-aggregate functions can be computed with no leader, even when restricted to simple connected static networks.
Proof.
Let $m_{1}$, $m_{2}$, …, $m_{k}$ be integers greater than $2$ with $\gcd(m_{1},m_{2},\dots,m_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $m_{1}$, $m_{2}$, …, $m_{k}$, respectively. For any positive integer $\alpha$, construct the static network $G_{\alpha}$ consisting of $\alpha$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $\alpha m_{i}$ processes in the $\alpha$ copies of $V_{i}$. Clearly, $G_{\alpha}$ is a simple connected static network.
Due to the fundamental theorem of history trees [24, Theorem 3.1], all the processes with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{\alpha}})=\psi(z_{i},\alpha\cdot\mu_{\lambda})$, regardless of $\alpha$. Hence, by definition, only frequency-based multi-aggregate functions can be computed in these networks. ∎
Proposition 5.2.
No algorithm can solve the Average Consensus problem in a $T$-union-connected leaderless network in less than $2Tn-O(T)$ rounds (with or without termination).
Proof.
According to [24, Theorem 5.2], the number of processes $n$ in a network with $\ell=1$ and $T=1$ cannot be determined in less than $2n-O(1)$ rounds (with or without termination). We can reduce this problem to Average Consensus with $\ell=0$ and $T=1$ as follows. In any given network with $\ell=T=1$, assign input $1$ to the leader and clear its leader flag; assign input 00 to all other processes. If the processes can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $T=1$ cannot be solved in less than $2n-O(1)$ rounds; this immediately generalizes to an arbitrary $T$ by Proposition 2.3. ∎
Proposition 5.3.
No algorithm can solve the leaderless Average Consensus problem with explicit termination if nothing is known about the size of the network, even when restricted to simple connected static networks.
Proof.
5.2 Networks with Leaders
Proposition 5.4.
No function other than the multi-aggregate functions can be computed (with or without termination), even when restricted to simple connected static networks with a known number of leaders.
Proof.
It is sufficient to construct a static network where all processes with the same input are indistinguishable at every round. Such is, for example, the complete graph $K_{n}$, where the output of a process can only depend on its input and the multiset of all processes’ inputs [24, Theorem 5.1]. This is the definition of a multi-aggregate function, and thus no other functions can be computed. ∎
Proposition 5.5.
No algorithm can compute the Counting function $F_{C}$ (with or without termination) with no knowledge about $\ell$, even when restricted to simple connected static networks with a known and arbitrarily small ratio $\ell/n$.
Proof.
Proposition 5.6.
For any $\ell\geq 1$, no algorithm can compute the Counting function $F_{C}$ (with or without termination) in all simple $T$-union-connected networks with $\ell$ leaders in less than $T(2n-\ell)-O(T)$ rounds.
Proof.
It was shown in [24, Theorem 5.2] that there is a family of simple $1$-union-connected networks $\mathcal{G}_{n}$, with $n\geq 1$, with the following properties. $\mathcal{G}_{n}$ has $\ell=1$ leader and $n$ processes in total; moreover, up to round $2n-O(1)$, the leaders of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have isomorphic views.
It is straightforward to see that the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n+1}$, which correspond to the leaders of $\mathcal{G}_{n-\ell+1}$ and $\mathcal{G}_{n-\ell+2}$ respectively, have isomorphic views up to round $2(n-\ell)-O(1)$. Since the view of $p_{1}$ is completely determined by the view of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the process $p_{1}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{1}$ in $\mathcal{G}^{\prime}_{n+1}$ have isomorphic views up to round $2n-\ell-O(1)$.
Thus, up to that round, the two processes must give an equal output, implying that they cannot both output the number of processes in their respective networks. It follows that the Counting function with $\ell\geq 1$ leaders and $T=1$ cannot be computed in less than $2n-\ell-O(1)$ rounds, which generalizes to an arbitrary $T$ by Proposition 2.3. ∎
6 Conclusions
We have shown that anonymous processes in disconnected dynamic networks can compute all the multi-aggregate functions and no other functions, provided that the network contains a known number of leaders $\ell\geq 1$. If there are no leaders or the number of leaders is unknown, the class of computable functions reduces to the frequency-based multi-aggregate functions. We have also identified the functions $F_{GC}$ and $F_{R}$ as the complete problems for each class. Notably, the network’s dynamic disconnectivity $T$ does not affect the computability of functions, but only makes computation slower.
Moreover, we gave efficient stabilizing and terminating algorithms for computing all the aforementioned functions. Some of our algorithms make assumptions on the processes’ a-priori knowledge about the network; we proved that such assumptions are actually necessary. All our algorithms have optimal linear running times in terms of $T$ and the size of the network $n$.
In one case, there is still a small gap in terms of the number of leaders $\ell$. Namely, for terminating computation with $\ell\geq 1$ leaders, we have a lower bound of $T(2n-\ell)-O(T)$ rounds (Proposition 5.6) and an upper bound of $(\ell^{2}+\ell+1)Tn$ rounds (Theorem 4.3). Although these bounds asymptotically match if the number of leaders $\ell$ is constant (which is a realistic assumption in most applications), optimizing them with respect to $\ell$ is left as an open problem.
Observe that our stabilizing algorithms use an unbounded amount of memory, as processes keep adding nodes to their view at every round. This can be avoided if the dynamic disconnectivity $T$ (as well as an upper bound on $n$, in case of a leaderless network) is known: In this case, processes can run the stabilizing and the terminating version of the relevant algorithm in parallel, and stop adding nodes to their views when the terminating algorithm halts. It is an open problem whether a stabilizing algorithm for $F_{GC}$ or $F_{R}$ can use a finite amount of memory with no knowledge of $T$.
Our algorithms require processes to send each other explicit representations of their history trees, which have cubic size in the worst case [24]. It would be interesting to develop algorithms that only send messages of logarithmic size, possibly with a trade-off in terms of running time. We are currently able to do so for leaderless networks and networks with a unique leader, but not for networks with more than one leader [25].
We also wonder if our results hold more generally for networks where communications are not necessarily synchronous. We conjecture that our algorithms can be generalized to networks where messages may be delayed by a bounded number of rounds or processes may be inactive for some rounds (provided that a “global fairness” condition is met).
APPENDIX
Appendix A The Subroutine ApproxCount and Its Correctness

### 05 Apr 2024 v6

@@ 5.2 Main Algorithm
The subroutine ApproxCount. We first introduce the subroutine ApproxCount, whose formal description and proof of correctness are postponed to Section 5.3. The purpose of ApproxCount is to compute an approximation $n^{\prime}$ of the total number of processes $n$ (or report various types of failure). It takes as input a view $\mathcal{V}$ of a process, the number of leaders $\ell$, and two integer parameters $s$ and $x$, representing the index of a level of $\mathcal{V}$ and the anonymity of a leader node in $L_{s}$, respectively.
@@
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\vartheta$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\vartheta$ does not exist, the procedure immediately returns the error code $n^{\prime}=-1$). We define the discrepancy $\delta$ as the ratio $x/a(\vartheta)$. Clearly, $\delta\leq\ell$. Note that, since $a(\vartheta)$ is not a-priori known by the process executing ApproxCount, then neither is $\delta$.
@@
Conditional anonymities. ApproxCount starts by assuming that the anonymity of $\vartheta$ is $x$, and makes deductions on other anonymities based on this assumption. Thus, we will distinguish between the actual anonymity of a node $a(v)$ and the conditional anonymity $a^{\prime}(v)=\delta a(v)$ that ApproxCount may compute under the initial assumption that $a^{\prime}(\vartheta)=x=\delta a(\vartheta)$.
@@
Overview of ApproxCount. The procedure ApproxCount scans the levels of $\mathcal{V}$ starting from $L_{s}$, making “guesses” on the conditional anonymities of nodes based on already known conditional anonymities. Generalizing some lemmas from [23], we develop a criterion to determine when a guess is correct. This yields more nodes with known conditional anonymities, and therefore more guesses (the details are in Section 5.3). As soon as it has obtained enough information, the procedure stops and returns $(n^{\prime},t)$, where $L_{t}$ is the level scanned thus far. If the information gathered satisfies certain criteria, then $n^{\prime}$ is an approximation of $n$. Otherwise, $n^{\prime}$ is an error code, as detailed below.
@@
Error codes. If $L_{s}$ contains no leader nodes, the procedure returns the error code $n^{\prime}=-1$. If, before gathering enough information on $n$, the procedure encounters a descendant of $\vartheta$ with more than one child in $\mathcal{V}$, it returns the error code $n^{\prime}=-2$. If it determines that the conditional anonymity of a node is not an integer, it returns the error code $n^{\prime}=-3$. Finally, if it determines that the sum $\ell^{\prime}$ of the conditional anonymities of the leader nodes is not $\ell$, it returns $n^{\prime}=-1$ if $\ell^{\prime}<\ell$ and $n^{\prime}=-3$ if $\ell^{\prime}>\ell$.
@@
Correctness of ApproxCount. The following lemma gives some conditions that guarantee that ApproxCount has the expected behavior; it also gives bounds on the number of rounds it takes for ApproxCount to produce an approximation $n^{\prime}$ of $n$, as well as a criterion to determine if $n^{\prime}=n$. The lemma’s proof is rather lengthy and technical, and is found in Section 5.3 (cf. Lemma 5.9).
@@
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let ϑsuperscriptitalic-$\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@
    If $x=a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}\neq-3$.
@@
    If $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$ and $a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}=n$.
@@
    If $t^{\prime}\geq s+(\ell+2)(n-1)$, then $s\leq t\leq s+(\ell+1)(n-1)$ and $n^{\prime}\neq-1$. Moreover, if $n^{\prime}=-2$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.∎
Our terminating algorithm assumes that all processes know the number of leaders $\ell\geq 1$ and the dynamic disconnectivity $\tau$. Again, this is justified by Proposition 6.2 and Proposition 2.3.
@@
There is an algorithm that computes $F_{IM}$ in all $\tau$-union-connected anonymous networks with $\ell\geq 1$ leaders and terminates in at most $\tau((\ell^{2}+\ell+1)(n-1)+1)$ rounds, assuming that $\ell$ and $\tau$ are known to all processes, but assuming no knowledge of $n$.
@@
Due to Proposition 2.4, since $\tau$ is known and appears as a factor in the claimed running time, we can assume that $\tau=1$ without loss of generality. Also, note that determining $n$ is enough to compute $F_{IM}$. Indeed, if a process determines $n$ at round $t^{\prime}$, it can wait until round $\max\{t^{\prime},\tau(2n-2)\}$ and run the algorithm in Theorem 4.7, which is guaranteed to give the correct output by that time.
In order to determine $n$ assuming that $\tau=1$, we let each process run the algorithm in Listing 2 with input $(\mathcal{V},\ell)$, where $\mathcal{V}$ is the view of the process at the current round $t^{\prime}$. We will prove that this algorithm returns a positive integer (as opposed to “Unknown”) within $(\ell^{2}+\ell+1)(n-1)+1$ rounds, and the returned number is indeed the correct size of the system $n$.
@@ 8   For $x:=\ell$ downto $1$
9      Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$      # see Listing 3 in Section 5.3
@@ 17   Assign $s:=t^{\ast}+1$
In order to compute an approximation of $n$, say in an interval of levels $\mathcal{L}_{i}$ starting at $L_{s}$, the algorithm goes through at most $\ell$ phases (Lines 8–16). The first phase begins by calling ApproxCount with starting level $L_{s}$ and $x=\ell$, i.e., the maximum possible value for the anonymity of a leader node (Line 9). Specifically, ApproxCount chooses a leader node in $\vartheta\in L_{s}$ and tries to estimate $n$ using as few levels as possible.
Let $(n^{\prime},t)$ be the pair of values returned by ApproxCount. If $n^{\prime}=-1$, this is evidence that $\mathcal{V}$ is still missing some relevant nodes, and therefore “Unknown” is immediately returned (Line 11). If $n^{\prime}=-2$, then a descendant of $\vartheta$ with multiple children in $\mathcal{V}$ was found, say at level $L_{t}$, before an approximation of $n$ could be determined. As this is an undesirable event, the algorithm moves $\mathcal{L}_{i}$ after $L_{t}$ and tries again to estimate $n$ (Line 12). If $n^{\prime}=-3$, then $x$ may not be the correct anonymity of the leader node $\vartheta$ (see the description of ApproxCount), and therefore the algorithm calls ApproxCount again with the same starting level $L_{s}$, but now with $x=\ell-1$. If $n^{\prime}=-3$ is returned again, then $x=\ell-2$ is tried, and so on. After all possible assignments down to $x=1$ have failed, the algorithm just moves $\mathcal{L}_{i}$ forward and tries again from $x=\ell$.
@@
Finally, when $\ell-b+1$ approximations of $n$ (all equal to $n^{\ast}$) have been found, a correctness check is performed: the algorithm takes the last level $L_{t^{\ast}}$ visited thus far; if the current round $t^{\prime}$ satisfies $t^{\prime}\geq t^{\ast}+n^{\ast}$, then $n^{\ast}=n$ is accepted as correct and returned; otherwise “Unknown” is returned (Lines 18–20).
@@
Correctness and running time. We will prove that, if the output of Listing 2 is not “Unknown”, then it is indeed the number of processes, i.e., $n^{\ast}=n$. Since the $\ell-b+1$ approximations of $n$ have been computed on disjoint intervals of levels, there is at least one such interval, say $\mathcal{L}_{j}$, where no leader node in the history tree has more than one child (because there can be at most $\ell$ leader branches). With the notation of Lemma 5.1, this implies that $a(\vartheta)=a(\vartheta^{\prime})$ whenever ApproxCount is called in $\mathcal{L}_{j}$. Also, since the option $x=\ell$ is tried first, the assumption $x\geq a(\vartheta)$ of Lemma 5.1 is initially satisfied. Note that ApproxCount cannot return $n^{\prime}=-1$ or $n^{\prime}=-2$, or else $\mathcal{L}_{j}$ would not yield any approximation of $n$. Moreover, by Lemma 5.1 (ii) and by the terminating condition (Line 19), if $n^{\prime}>0$ while $x\geq a(\vartheta)$, then $n^{\ast}=n^{\prime}=n$. On the other hand, due to Lemma 5.1 (i), by the time $x=a(\vartheta)$ we necessarily have $n^{\prime}\neq-3$ and therefore $n^{\prime}>0$.
@@
It remains to prove that Listing 2 actually gives an output other than “Unknown” within the claimed number of rounds; it suffices to show that it does so if it is executed at round $t^{\prime}=(\ell^{2}+\ell+1)(n-1)+1$. By Corollary 3.4, all nodes in the first $t^{\prime}-n+1=\ell(\ell+1)(n-1)+1$ levels of the history tree are contained in the view $\mathcal{V}$ at round $t^{\prime}$. It is straightforward to prove by induction that the assumption $t^{\prime}\geq s+(\ell+2)(n-1)$ of Lemma 5.1 (iii) holds every time ApproxCount is invoked. Indeed, as long as this condition is satisfied, Lemma 5.1 (iii) implies that $n^{\prime}\neq-1$, and so “Unknown” is not returned at Line 11. Also, reasoning as in the previous paragraph, we infer that $n^{\prime}\neq-3$ by the time $x=a(\vartheta)$. Thus, within $\Delta=(\ell+1)(n-1)$ levels, either a branching leader node is found ($n^{\prime}=-2$) or a new approximation of $n$ is computed ($n^{\prime}>0$). Every time either event occurs, $s$ is increased by at most $\Delta$ at Line 17. Thus, after $\ell-1$ (or fewer) updates of $s$, we have $s\leq(\ell-1)\Delta=t^{\prime}-(\ell+2)(n-1)-1$. Hence the condition of Lemma 5.1 (iii) holds again, and the induction goes through for at least $\ell$ steps.
Observe that, since there can be at most $\ell-1$ branching leader nodes in $\mathcal{V}$, at least one approximation $n^{\prime}>0$ of $n$ is computed within the $\ell$th iteration of the loop at Lines 6–17. This occurs within $t^{\ast}\leq\ell\Delta=t^{\prime}-n$ levels. Because all nodes in these levels must appear in $\mathcal{V}$, the condition $a(\vartheta)=a(\vartheta^{\prime})$ is satisfied in all intervals $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, …, $\mathcal{L}_{\ell-b+1}$. Due to Lemma 5.1 (ii), we conclude that all such intervals must yield the correct approximation $n^{\prime}=n$. So, every time Line 15 is executed, we have $n^{\ast}=n^{\prime}$, and the algorithm does not return “Unknown”. Finally, when Line 19 is reached, we have $t^{\ast}\leq t^{\prime}-n=t^{\prime}-n^{\ast}$, and therefore “Unknown” is not returned. Thus, an output other than “Unknown” is returned within the desired number of rounds. ∎
5.3 Subroutine ApproxCount
We will now define the subroutine $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ introduced in Section 5.2 and invoked in Listing 2. We will also give a proof to the technical Lemma 5.1 (restated as Lemma 5.9).